# disposition

### v2

@@ Appendix C Time Series Asymptotics
Given that $\overline{\mathbf{W}}_{n,i}$ is $\alpha$-mixing of size $-2r/(r-2)$ for some $r>2$, we have $\widehat{\Omega}_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\Omega$.
@@ Appendix D Simulations
We analyze the finite sample performance of the confidence intervals described in Section 6.2 based on the standard asymptotics in Propositions 6.46.6 as well as the Fisher transformations described in Section 6.3, Appendix B and Propositions B.1B.3. We simulate the pairs of binary random variables $(X_{i},Y_{i})$ independently for different $i=1,\dots,n$ and use $n\in\{100;500;2{,}000\}$. The joint behavior of $(X_{i},Y_{i})$ is characterized by the probabilities $p=\mathbb{P}(X_{i}=1)$, $q=\mathbb{P}(Y_{i}=1)$ and $r=\mathbb{P}(X_{i}=1,\,Y_{i}=1)$. In order to cover the different settings in Proposition 6.5, we use $(p,q)\in\big{\{}(0.05,0.5),\,(0.1,0.9),\,(0.2,0.2),\,(0.3,0.7),\,(0.4,0.8)\big{\}}$. For each of these combinations of $p$ and $q$, we choose $39$ equally spaced values for $r$ between $\max(p+q-1,0)+1/20$ and $\min(p,q)-1/20$. In the following, we only report results for the settings where the empirical boundary-conditions $10<\widehat{{p}}_{n},\widehat{{q}}_{n}<$ 1 and $\max(\widehat{{p}}_{n}+\widehat{{q}}_{n}-1,0)<\widehat{{r}}_{n}<\min(\widehat{{p}}_{n},\widehat{{q}}_{n})$ are satisfied in at least $50\%$ of the simulation runs.
@@
Figure 7: Empirical confidence interval coverage rates with nominal level $0.91-\alpha=$ 0.9 for $M=1{,}000$ simulation replications for the measures ${Q}$ (top), ${C}$ (middle) and $\phi$ (bottom) for the sample sizes $n\in\{100;500;2{,}000\}$ and the probabilities $p$ and $q$ given in the column headings. The $x$-axis displays the true values for $\{{Q},{C},\phi\}$ that arise for the sequence of joint probabilities $r$. The respective Fréchet–Hoeffding bounds are displayed as vertical dashed lines. Standard confidence intervals are given in red and confidence intervals based on the Fisher transformation in blue.
@@
Figure 7 shows the coverage rates of the confidence intervals that are constructed based on the standard and the Fisher-transformed asymptotic distributions for ${Q}$, ${C}$ and $\phi$ in the respective panels. For ${Q}$, we obtain very accurate coverage rates that converge to the nominal level of $90\%$. While the standard approach displays inaccuracies close to the boundaries, the Fisher transformation expectedly improves the accuracy in these regions.
@@
For ${C}$, we observe mostly conservative coverage rates that are caused by the multiple testing corrections in the construction of the confidence intervals. The Fisher transformation again improves the accuracy close to the boundaries. The notable downside peak in the coverage rates close to the value of ${C}=0$ can be explained by the piecewise normalization of ${C}$. In particular, for a true $\sigma>0$, the last two terms in (19) involving the quantity $\mathds{1}\{\widehat{\sigma}_{n}<0\}$ in the proof of Proposition 6.5 are shown to be $o_{\mathbb{P}}(1)$ and hence do not contribute to the asymptotic distribution, but deliver non-negligible contributions to the finite sample distributions.
@@
For $\phi$, we again see the drastic effect of the non-attainability through the theoretical boundaries depicted with the vertical dashed lines. While the coverage rates of the standard method are very accurate in the center, we observe an over-coverage closer to the boundaries that cannot be improved with the Fisher transformation. As the Fisher transformation leaves the central part [0.5, 0.5]0.50.5[-0.5,\,0.5][ - 0.5 , 0.5 ] almost unchanged, it is ineffective when the theoretical boundaries of $\phi$ are far from $-1$ and $1$, once again illustrating a fundamental deficiency of the phi coefficient.
@@
Figure 8: Empirical violation rates of confidence intervals with nominal level $0.91-\alpha=$ 0.9. The “lower” (“upper”) violation rate reports the frequency of true values being below (above) the estimated confidence intervals. We show results for $M=1{,}000$ simulation replications for the measures ${Q}$ (top), ${C}$ (middle) and $\phi$ (bottom) for the sample sizes $n\in\{100;500;2{,}000\}$ and the probabilities $p$ and $q$ given in the column headings. The $x$-axis displays the true values for $\{{Q},{C},\phi\}$ that arise for the sequence of joint probabilities $r$. The respective Fréchet–Hoeffding bounds are displayed as vertical dashed lines. Standard confidence intervals are given in red and confidence intervals based on the Fisher transformation in blue.
@@
Figure 8 refines the analysis by plotting the lower (upper) “violation rates”, which are defined as the frequency of population dependence measures being below (above) the estimated confidence intervals. For ${Q}$ in the upper panel, we especially see the asymmetry of the standard asymptotics that can be explained by approximating the skewed empirical distribution with the symmetric Gaussian limit; see in particular Figure 6. The Fisher transformation fixes this asymmetry and delivers confidence intervals that have approximately the same mass to either side. While similar effects can be observed for ${C}$, the Fisher transformation is again relatively ineffective for $\phi$.
@@
Basic $\mathcal{H}_{c}=[\mathcal{H}_{=}\cup\mathcal{H}_{\not=}]$ $\max[2p_{=},p_{\not=}]$
@@
For $c>0$: ==$\mathcal{H}_{=}=\{C=c,p=q\},\;\mathcal{H}_{p,q}=\{p=q\},\;\mathcal{H}_{\sigma}=\{\sigma\geq 0\}$
For $c<0$: ==$\mathcal{H}_{=}=\{C=c,p=1-q\},\;\mathcal{H}_{\not=}=\{C=c,p\not=1-q\},\;\mathcal{H}_{p,q}=\{p=1-q\},\;\mathcal{H}_{\sigma}=\{\sigma\leq 0\}$
Table 5: Overview of the different theoretically valid combination methods of elementary hypotheses and corresponding $p$-values used for the construction of confidence intervals for ${C}$ in Figure 9.
Figure 9: Empirical coverage rates (upper panel) and average confidence interval length (lower panel) with level $0.91-\alpha=$ 0.9 based on the different construction principles given in Table 5 for ${C}$ using the Fisher transformation. The $x$-axes display the true values of ${C}$ that arise for the sequence of joint probabilities $r$.
We continue to illustrate the benefits of the combination of elementary hypotheses in the construction of tests for $\mathcal{H}_{c}=\{C=c\}$ for any $c\not=0$ as discussed in Section 6.2. For this, we consider simplified methods that omit one/both tests for $\mathcal{H}_{p,q}$ and $\mathcal{H}_{\sigma}$ respectively as illustrated in detail in Table 5. Figure 9 displays the empirical coverage rates and average lengths of confidence intervals for ${C}$ based on these different construction principles for a fixed sample size of $n=500$ and by using the Fisher transformation. We see that all confidence intervals have satisfactory coverage rates that are often conservative due to the multiple testing corrections. However, the lengths of the different methods differ substantially, whereas the “full” method we propose achieves overall very narrow intervals. While it is often not the most narrow one as its inherent multiple testing corrections result in conservative coverage rates, it is in all settings close to the most narrow method. Which method is the most narrow one however depends on practically unknown population quantities, and hence cannot be known in practice. In contrast, omitting the respective tests for $\mathcal{H}_{p,q}$ and $\mathcal{H}_{\sigma}$ results in drastic increases of the length in certain situations. Overall, the “full” method is relatively narrow (powerful) in all settings and trades some length in well-behaved settings in order to insure against very unfortunate cases.
@@ E.1 Proofs for the Results in Sections 25 and Appendix A
[…]
@@
The equivalence of (i) and (ii) is established by checking under which conditions the bounds from (3) become 1 and $-1$.
@@
Denote the Fréchet–Hoeffding upper and lower bound of ${Cov}(A,B)$ from (2) by
@@
By part (i) of Proposition 2.4 and Lemma 2.5 (iii) it holds that ${FH}^{+}(A,\overline{B})={Cov}(A,\overline{B})$ and ${FH}^{-}(A,B)={Cov}(A,B)$ if $A$ and $B$ are perfectly negatively dependent. By the second part of property (E) of ${Cov}(A,B)$ (as established in Proposition 4.2) the claim follows. ∎
@@
(B): As the normalization terms are positive it holds that ${C}(A,B)=0$ if and only if ${Cov}(A,B)=0$ and since ${Cov}(A,B)$ fulfills property (B) by Proposition 4.2, ${C}(A,B)$ does so as well.
(D): For fixed marginal probabilities ${C}(A,B)$ is a strictly monotonic function of ${Cov}(A,B)$ and thus fulfills property (D) because ${Cov}(A,B)$ does.
(E): The normalization terms and ${Cov}(A,B)$ are symmetric and thus ${C}(A,B)$ is as well, ${C}(A,B)=C(B,A)$. For the second symmetry property, we make a case distinction: For ${Cov}(A,B)=0$, the analogous property of covariance implies $0={Cov}(A,\overline{B})=-{Cov}(A,B)$, which establishes the claim (as the normalization is nonzero). For ${Cov}(A,B)>0$, we have by the analogous property of ${Cov}(A,B)$ and by Lemma E.1 that ${C}(A,\overline{B})=\frac{{Cov}(A,\overline{B})}{{FH}^{+}(A,\overline{B})}=\frac{{Cov}(A,B)}{{FH}^{-}(A,B)}=-{C}(A,B)$. The proof works analogously for ${Cov}(A,B)<0$. ∎
@@ E.2 Proofs for the Asymptotic Results in Section 6
As $\widehat{{p}}_{n}$ and $\widehat{{q}}_{n}$ are consistent estimators for $p$ and $q$ and as Assumption 6.1 is in place, we have that ($\mathbb{P}\big{(}\{\widehat{{p}}_{n}\in\{0,1\}\}\cup\{\widehat{{q}}_{n}\in\{0,1\}\}\big{)}\to 0$ such that $\widehat{{Q}}_{n}$ is well-defined asymptotically with probability one. Formally, for any set $D\in\mathcal{F}$, we have that
($\mathbb{P}\big{(}\{\sqrt{n}\big{(}\widehat{{Q}}_{n}-{Q}\big{)}\in D\}\big{)}&=\mathbb{P}\big{(}\{\sqrt{n}\big{(}\widehat{{Q}}_{n}-{Q}\big{)}\in D\}\cap\{0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1\}\big{)}\\ &\qquad+\mathbb{P}\big{(}\{\sqrt{n}\big{(}\widehat{{Q}}_{n}-{Q}\big{)}\in D\}\cap\{\widehat{{p}}_{n}\in\{0,1\}\text{ or }\widehat{{q}}_{n}\in\{0,1\}\}\big{)}\\ &=\mathbb{P}\big{(}\{\sqrt{n}\big{(}\widehat{{Q}}_{n}-{Q}\big{)}\in D\}\cap\{0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1\}\big{)}+o_{\mathbb{P}}(1),\end{aligned}$ (15)
@@
which is differentiable at $(p,q,r)$ as $(q-r)(p-r)>0$ and $r(1-p-q+r)>0$ by Assumption 6.1. The function $g$ has the (non-zero) Jacobian matrix of continuous partial derivatives
@@
$\bullet$ In the zero case, let $\sigma={Cov}(A,B)=0$ such that ${C}=0$. We get from (18) that
@@
[…]
$\bullet$ In the positive case, let $\sigma={Cov}(A,B)>0$ such that ${C}>0$. Then,
[…]
@@
$\bullet$ In the negative case, let $\sigma={Cov}(A,B)<0$ such that ${C}<0$. With arguments as above,
@@
Second, for the case $p=1-q$, we again use the random vector $\mathbf{V}_{n}=\sqrt{n}\big{(}\widehat{{p}}_{n}-p,\,\widehat{{q}}_{n}-q,\,\widehat{{p}}_{n}\widehat{{q}}_{n}-pq,\,\widehat{\sigma}_{n}-\sigma\big{)}$ together with its (degenerate) limiting distribution $\mathbf{V}=J_{f}\Omega^{1/2}\mathbf{Z}$, where $\mathbf{Z}\sim\mathcal{N}(0,I_{3})$. We then get that

### v3

@@ Appendix C Time Series Asymptotics
Given that $\overline{\mathbf{W}}_{n,i}$ is $\alpha$-mixing of size 2r/(r2)-2r/(r-2) for some $r>2$, we have $\widehat{\Omega}_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\Omega$.
@@ Appendix D Simulations
We analyze the finite sample performance of the confidence intervals described in Section 6.2 based on the standard asymptotics in Propositions 6.46.6 as well as the Fisher transformations described in Section 6.3, Appendix B and Propositions B.1B.3. We simulate the pairs of binary random variables $(X_{i},Y_{i})$ independently for different $i=1,\dots,n$ and use $n\in\{100;500;2{,}000\}$. The joint behavior of $(X_{i},Y_{i})$ is characterized by the probabilities $p=\mathbb{P}(X_{i}=1)$, $q=\mathbb{P}(Y_{i}=1)$ and $r=\mathbb{P}(X_{i}=1,\,Y_{i}=1)$. In order to cover the different settings in Proposition 6.5, we use $(p,q)\in\big\{(0.05,0.5),\,(0.1,0.9),\,(0.2,0.2),\,(0.3,0.7),\,(0.4,0.8)\big\}$. For each of these combinations of $p$ and $q$, we choose $39$ equally spaced values for $r$ between $\max(p+q-1,0)+1/20$ and $\min(p,q)-1/20$. In the following, we only report results for the settings where the empirical boundary-conditions $0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1$ and $\max(\widehat{{p}}_{n}+\widehat{{q}}_{n}-1,0)<\widehat{{r}}_{n}<\min(\widehat{{p}}_{n},\widehat{{q}}_{n})$ are satisfied in at least $50\%$ of the simulation runs.
@@
Figure 7: Empirical confidence interval coverage rates with nominal level $1-\alpha=0.9$ for M=1, [^0] M=1{,}000 simulation replications for the measures Q{Q} (top), C{C} (middle) and $\phi$ (bottom) for the sample sizes $n\in\{100;500;2{,}000\}$ and the probabilities $p$ and $q$ given in the column headings. The $x$-axis displays the true values for $\{{Q},{C},\phi\}$ that arise for the sequence of joint probabilities $r$. The respective Fréchet–Hoeffding bounds are displayed as vertical dashed lines. Standard confidence intervals are given in red and confidence intervals based on the Fisher transformation in blue.
@@
Figure 7 shows the coverage rates of the confidence intervals that are constructed based on the standard and the Fisher-transformed asymptotic distributions for Q{Q}, C{C} and $\phi$ in the respective panels. For Q{Q}, we obtain very accurate coverage rates that converge to the nominal level of $90\%$. While the standard approach displays inaccuracies close to the boundaries, the Fisher transformation expectedly improves the accuracy in these regions.
@@
For C{C}, we observe mostly conservative coverage rates that are caused by the multiple testing corrections in the construction of the confidence intervals. The Fisher transformation again improves the accuracy close to the boundaries. The notable downside peak in the coverage rates close to the value of C=0{C}=0 can be explained by the piecewise normalization of C{C}. In particular, for a true $\sigma>0$, the last two terms in (19) involving the quantity $\mathds{1}\{\widehat{\sigma}_{n}<0\}$ in the proof of Proposition 6.5 are shown to be $o_{\mathbb{P}}(1)$ and hence do not contribute to the asymptotic distribution, but deliver non-negligible contributions to the finite sample distributions.
@@
For $\phi$, we again see the drastic effect of the non-attainability through the theoretical boundaries depicted with the vertical dashed lines. While the coverage rates of the standard method are very accurate in the center, we observe an over-coverage closer to the boundaries that cannot be improved with the Fisher transformation. As the Fisher transformation leaves the central part [0.5, 0.5][-0.5,\,0.5] almost unchanged, it is ineffective when the theoretical boundaries of $\phi$ are far from 1-1 and $1$, once again illustrating a fundamental deficiency of the phi coefficient.
@@
Figure 8: Empirical violation rates of confidence intervals with nominal level $1-\alpha=0.9$. The “lower” (“upper”) violation rate reports the frequency of true values being below (above) the estimated confidence intervals. We show results for M=1, [^0] M=1{,}000 simulation replications for the measures Q{Q} (top), C{C} (middle) and $\phi$ (bottom) for the sample sizes $n\in\{100;500;2{,}000\}$ and the probabilities $p$ and $q$ given in the column headings. The $x$-axis displays the true values for $\{{Q},{C},\phi\}$ that arise for the sequence of joint probabilities $r$. The respective Fréchet–Hoeffding bounds are displayed as vertical dashed lines. Standard confidence intervals are given in red and confidence intervals based on the Fisher transformation in blue.
@@
Figure 8 refines the analysis by plotting the lower (upper) “violation rates”, which are defined as the frequency of population dependence measures being below (above) the estimated confidence intervals. For Q{Q} in the upper panel, we especially see the asymmetry of the standard asymptotics that can be explained by approximating the skewed empirical distribution with the symmetric Gaussian limit; see in particular Figure 6. The Fisher transformation fixes this asymmetry and delivers confidence intervals that have approximately the same mass to either side. While similar effects can be observed for C{C}, the Fisher transformation is again relatively ineffective for $\phi$.
@@
Basic $\mathcal{H}_{c}=[\mathcal{H}_{=}\cup\mathcal{H}_{\not=}]$ $\max[p_{=},p_{\not=}]$
@@
For $c>0$: $\mathcal{H}_{=}=\{C=c,p=q\},\;\mathcal{H}_{\not=}=\{C=c,p\not=q\},\;\mathcal{H}_{p,q}=\{p=q\},\;\mathcal{H}_{\sigma}=\{\sigma\geq 0\}$
For $c<0$: $\mathcal{H}_{=}=\{C=c,p=1-q\},\;\mathcal{H}_{\not=}=\{C=c,p\not=1-q\},\;\mathcal{H}_{p,q}=\{p=1-q\},\;\mathcal{H}_{\sigma}=\{\sigma\leq 0\}$
Table 5: Overview of the different theoretically valid combination methods of elementary hypotheses and corresponding $p$-values used for the construction of confidence intervals for C{C} in Figure 9.
Figure 9: Empirical coverage rates (upper panel) and average confidence interval length (lower panel) with level $1-\alpha=0.9$ based on the different construction principles given in Table 5 for C{C} using the Fisher transformation. The $x$-axes display the true values of C{C} that arise for the sequence of joint probabilities $r$.
We continue to illustrate the benefits of the combination of elementary hypotheses in the construction of tests for $\mathcal{H}_{c}=\{C=c\}$ for any $c\not=0$ as discussed in Section 6.2. For this, we consider simplified methods that omit one/both tests for $\mathcal{H}_{p,q}$ and $\mathcal{H}_{\sigma}$ respectively as illustrated in detail in Table 5. Figure 9 displays the empirical coverage rates and average lengths of confidence intervals for C{C} based on these different construction principles for a fixed sample size of $n=500$ and by using the Fisher transformation. We see that all confidence intervals have satisfactory coverage rates that are often conservative due to the multiple testing corrections. However, the lengths of the different methods differ substantially, whereas the “full” method we propose achieves overall very narrow intervals. While it is often not the most narrow one as its inherent multiple testing corrections result in conservative coverage rates, it is in all settings close to the most narrow method. Which method is the most narrow one however depends on practically unknown population quantities, and hence cannot be known in practice. In contrast, omitting the respective tests for $\mathcal{H}_{p,q}$ and $\mathcal{H}_{\sigma}$ results in drastic increases of the length in certain situations. Overall, the “full” method is relatively narrow (powerful) in all settings and trades some length in well-behaved settings in order to insure against very unfortunate cases.
@@ E.1 Proofs for the Results in Sections 25 and Appendix A
AA and $B$ being positively (negatively) dependent is equivalent to them being stronger positively (negatively) dependent than two independent events $A^{*}$ and $B^{*}$ with the same marginal probabilities. By (B) it holds that $\delta(A^{*},B^{*})=0$. Thus, by (D), $\delta(A,B)\geq(\leq)\ 0$ is equivalent to positive (negative) dependence of $A$ and $B$. ∎
@@
The equivalence of (i) and (ii) is established by checking under which conditions the bounds from (3) become 1 and 1-1.
@@
Denote the Fréchet–Hoeffding upper and lower bound of Cov(A,B){Cov}(A,B) from (2) by
@@
By part (i) of Proposition 2.4 and Lemma 2.5 (iii) it holds that ${FH}^{+}(A,\overline{B})={Cov}(A,\overline{B})$ and ${FH}^{-}(A,B)={Cov}(A,B)$ if $A$ and $B$ are perfectly negatively dependent. By the second part of property (E) of Cov(A,B){Cov}(A,B) (as established in Proposition 4.2) the claim follows. ∎
@@
(B): As the normalization terms are positive it holds that C(A,B)=0{C}(A,B)=0 if and only if Cov(A,B)=0{Cov}(A,B)=0 and since Cov(A,B){Cov}(A,B) fulfills property (B) by Proposition 4.2, C(A,B){C}(A,B) does so as well.
(D): For fixed marginal probabilities C(A,B){C}(A,B) is a strictly monotonic function of Cov(A,B){Cov}(A,B) and thus fulfills property (D) because Cov(A,B){Cov}(A,B) does.
(E): The normalization terms and Cov(A,B){Cov}(A,B) are symmetric and thus C(A,B){C}(A,B) is as well, C(A,B)=C(B,A){C}(A,B)=C(B,A). For the second symmetry property, we make a case distinction: For Cov(A,B)=0{Cov}(A,B)=0, the analogous property of covariance implies $0={Cov}(A,\overline{B})=-{Cov}(A,B)$, which establishes the claim (as the normalization is nonzero). For Cov(A,B)>0{Cov}(A,B)>0, we have by the analogous property of Cov(A,B){Cov}(A,B) and by Lemma E.1 that ${C}(A,\overline{B})=\frac{{Cov}(A,\overline{B})}{{FH}^{+}(A,\overline{B})}=\frac{{Cov}(A,B)}{{FH}^{-}(A,B)}=-{C}(A,B)$. The proof works analogously for Cov(A,B)<0{Cov}(A,B)<0. ∎
@@ E.2 Proofs for the Asymptotic Results in Section 6
As $\widehat{{p}}_{n}$ and $\widehat{{q}}_{n}$ are consistent estimators for $p$ and $q$ and as Assumption 6.1 is in place, we have that $\mathbb{P}\big(\{\widehat{{p}}_{n}\in\{0,1\}\}\cup\{\widehat{{q}}_{n}\in\{0,1\}\}\big)\to 0$ such that $\widehat{{Q}}_{n}$ is well-defined asymptotically with probability one. Formally, for any set $D\in\mathcal{F}$, we have that
$\begin{aligned} \mathbb{P}\big(\{\sqrt{n}\big(\widehat{{Q}}_{n}-{Q}\big)\in D\}\big)&=\mathbb{P}\big(\{\sqrt{n}\big(\widehat{{Q}}_{n}-{Q}\big)\in D\}\cap\{0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1\}\big)\\ &\qquad+\mathbb{P}\big(\{\sqrt{n}\big(\widehat{{Q}}_{n}-{Q}\big)\in D\}\cap\{\widehat{{p}}_{n}\in\{0,1\}\text{ or }\widehat{{q}}_{n}\in\{0,1\}\}\big)\\ &=\mathbb{P}\big(\{\sqrt{n}\big(\widehat{{Q}}_{n}-{Q}\big)\in D\}\cap\{0<\widehat{{p}}_{n},\widehat{{q}}_{n}<1\}\big)+o_{\mathbb{P}}(1),\end{aligned}$ (15)
@@
which is differentiable at $(p,q,r)$ as (qr)(pr)>0(q-r)(p-r)>0 and r(1pq+r)>0r(1-p-q+r)>0 by Assumption 6.1. The function $g$ has the (non-zero) Jacobian matrix of continuous partial derivatives
@@
$\bullet$ In the zero case, let $\sigma={Cov}(A,B)=0$ such that C=0{C}=0. We get from (18) that
@@
$\big(\Delta^{\top}\Omega\Delta\big)^{-1/2}\sqrt{n}\big(\widehat{{C}}_{n}-{C}\big)\stackrel{{\scriptstyle d}}{{\to}}\frac{Z\,\mathds{1}\{Z>0\}}{m^{+}}+\frac{Z\,\mathds{1}\{Z<0\}}{m^{-}}.$
$\bullet$ In the positive case, let $\sigma={Cov}(A,B)>0$ such that C>0{C}>0. Then,
$\begin{aligned} \sqrt{n}\big(\widehat{{C}}_{n}-{C}\big)&=\sqrt{n}\left(\frac{\widehat{\sigma}_{n}}{\widehat{m}_{n}^{+}}\mathds{1}\{\widehat{\sigma}_{n}>0\}-\frac{\sigma}{m^{+}}\right)+\sqrt{n}\frac{\widehat{\sigma}_{n}}{\widehat{m}_{n}^{-}}\mathds{1}\{\widehat{\sigma}_{n}<0\}\\ &=\sqrt{n}\frac{\widehat{\sigma}_{n}\mathds{1}\{\widehat{\sigma}_{n}>0\}-\sigma}{\widehat{m}_{n}^{+}}+\sqrt{n}\left(\frac{\sigma}{\widehat{m}_{n}^{+}}-\frac{\sigma}{m^{+}}\right)+\sqrt{n}\frac{\widehat{\sigma}_{n}}{\widehat{m}_{n}^{-}}\mathds{1}\{\widehat{\sigma}_{n}<0\}\\ &=\sqrt{n}\frac{\widehat{\sigma}_{n}-\sigma}{\widehat{m}_{n}^{+}}+\sqrt{n}\left(\frac{\sigma(m^{+}-\widehat{m}_{n}^{+})}{m^{+}\widehat{m}_{n}^{+}}\right)-\sqrt{n}\frac{\widehat{\sigma}_{n}}{\widehat{m}_{n}^{+}}\mathds{1}\{\widehat{\sigma}_{n}\leq 0\}+\sqrt{n}\frac{\widehat{\sigma}_{n}}{\widehat{m}_{n}^{-}}\mathds{1}\{\widehat{\sigma}_{n}<0\}.\end{aligned}$ (19)
@@
$\bullet$ In the negative case, let $\sigma={Cov}(A,B)<0$ such that C<0{C}<0. With arguments as above,
@@
Second, for the case p=1qp=1-q, we again use the random vector $\mathbf{V}_{n}=\sqrt{n}\big(\widehat{{p}}_{n}-p,\,\widehat{{q}}_{n}-q,\,\widehat{{p}}_{n}\widehat{{q}}_{n}-pq,\,\widehat{\sigma}_{n}-\sigma\big)$ together with its (degenerate) limiting distribution $\mathbf{V}=J_{f}\Omega^{1/2}\mathbf{Z}$, where $\mathbf{Z}\sim\mathcal{N}(0,I_{3})$. We then get that